OpenStudy (anonymous):

To the nearest hundredth, what is the area of a circle with a circumference of 42 units?

OpenStudy (anonymous):

we need to find the radius to find the area :\[C = 2\pi r\]lets plug in all the information for have\[42 = 2\pi r\]\[\frac{42}{2\pi } = r\]now plug in our new found radius into the area formula to get the answer \[A = \pi r^2\]

OpenStudy (anonymous):

ohh so it would be 42/6.28 :/

OpenStudy (anonymous):

if you are using 3.14 as pi then it would be that :)

OpenStudy (anonymous):

then i divide that right :/

OpenStudy (anonymous):

well what you get i will check you :)

OpenStudy (anonymous):

6.69 :/

OpenStudy (anonymous):

yep i got round there too so now that we know the radius we need to plug in the r for the area of a cirlce which is \[A = \pi (6.69)^2\]

OpenStudy (anonymous):

wait thats not my answer :/

OpenStudy (anonymous):

ohhh nv i got it lol so my answer would be 140.53 :/

OpenStudy (anonymous):

yep! :)

OpenStudy (anonymous):

ohhh i c thank u :D

OpenStudy (anonymous):

no problem :)